# DiffusionGPT[Uncaptioned image]: LLM-Driven Text-to-Image Generation System

### v1

@@ Abstract
Diffusion models have opened up new avenues for the field of image generation, resulting in the proliferation of high-quality models shared on open-source platforms. However, a major challenge persists in current text-to-image systems are often unable to handle diverse inputs, or are limited to single model results. Current unified attempts often fall into two orthogonal aspects: i) parse Diverse Prompts in input stage; ii) activate expert model to output. To combine the best of both worlds, we propose DiffusionGPT, which leverages Large Language Models (LLM) to offer a unified generation system capable of seamlessly accommodating various types of prompts and integrating domain-expert models. DiffusionGPT constructs domain-specific Trees for various generative models based on prior knowledge. When provided with an input, the LLM parses the prompt and employs the Trees-of-Thought to guide the selection of an appropriate model, thereby relaxing input constraints and ensuring exceptional performance across diverse domains. Moreover, we introduce Advantage Databases, where the Tree-of-Thought is enriched with human feedback, aligning the model selection process with human preferences. Through extensive experiments and comparisons, we demonstrate the effectiveness of DiffusionGPT, showcasing its potential for pushing the boundaries of image synthesis in diverse domains.
@@
Figure 1: We propose a unified generation system DiffusionGPT, which leverages Large Language Models (LLM) to seamlessly accommodating various types of prompts input and integrating domain-expert models for output. Our system is capable of parsing diverse forms of inputs, including Prompt-based, Instruction-based, Inspiration-based, and Hypothesis-based input types. It exhibits the ability to generate outputs of superior quality.
[…]
1 Introduction
@@
Recent years have witnessed the prevalence of diffusion models[6] in image generation tasks, revolutionizing image editing, stylization, and other related tasks. Both DALLE-2[14] and Imagen[17] are extremely good at generating images from text prompts. However, their non-open source nature has hindered widespread popularization and corresponding ecological development. The first open-source text-to-image diffusion model, known as Stable Diffusion (SD)[16], which has rapidly gained popularity and widespread usage. Various techniques tailored for SD, such as Controlnet[27], Lora, further paved the way for the development of SD and foster its integration into various applications. SDXL[10] is the latest image generation model tailored to deliver exceptional photorealistic outputs with intricate details and artistic composition. Moreover, the impact of SD extends beyond technical aspects. Community platforms such as Civitai, WebUI, and LibLibAI have emerged as vibrant hubs for discussions and collaborations among designers and creators. The capability evolution of Stable Diffusion, the advancements in SD-customized technologies, and the thriving community platforms, has created an exciting and fertile ground for further advancements in image generation.
@@
Despite making significant strides, current stable diffusion models face two key challenges when applied to realistic scenarios:
  • Model Limitation: While stable diffusion (such as SD1.5) demonstrate adaptability to a wide range of prompts, they exhibit poor performance in specific domains. Conversely, domain-specific models (such as SD1.5+Lora) excel in producing extreme generation performance within specific sub-fields but lack versatility.
@@
  • Prompt Constraint: During the training of stable diffusion, text information typically consists of descriptive statements, such as captions. However, when utilizing these models, we encounter various prompt types, including instructions and inspirations. The current generation models struggle to achieve optimal generation performance for these diverse prompt types.
@@
The existing mix-matches between stable diffusion models and realistic applications often manifest as limited results, poor generalization, and increased difficulty in their practical implementation. A series of research works have explored to address these challenges and bridge the gap. While SDXL has made notable advancements in improving specific-domain performance, achieving the ultimate performance in this regard remains elusive. Other approaches involve incorporating prompt engineering techniques or fixed prompt templates to enhance the quality of input prompts and improve the overall generation output. While these approaches have shown varying degrees of success in mitigating the aforementioned challenges, they do not provide a comprehensive solution. This leads us to pose a fundamental question: Can we create a unified framework to unleash prompt constraint and activate corresponding domain expert model ?
@@
In order to address the aforementioned question, we propose DiffusionGPT, which leverages Large Language Model (LLM) to offer a one-for-all generation system that seamlessly integrates superior generative models and effectively parses diverse prompts. DiffusionGPT constructs a Tree-of-Thought (ToT) structure, which encompasses various generative models based on prior knowledge and human feedback. When presented with an input prompt, the LLM first parses the prompt and then guides the ToT to identify the most suitable model for generating the desired output. Furthermore, we introduce Advantage Databases, where the Tree-of-Thought is enriched with valuable human feedback, aligning the LLM’s model selection process with human preferences.
@@
  • New Insight: DiffusionGPT employs Large Language Model (LLM) to drive the entire text-to-image generation system. The LLM acts as the cognitive engine, processing diverse inputs and facilitating expert selection for output.
@@
  • All-in-one System: DiffusionGPT provides a versatile and professional solution by being compatible with a wide range of diffusion models. Unlike existing approaches that are limited to descriptive prompts, our framework can handle various prompt types, expanding its applicability.
@@
  • Efficiency and Pioneering: DiffusionGPT stands out for its training-free nature, allowing for easy integration as a plug-and-play solution. Through the incorporation of the Tree-of-Thought (ToT) and human feedback, our system achieves higher accuracy and pioneers a flexible process for aggregation of more experts.
@@
  • High Effectiveness: DiffusionGPT outperforms traditional stable diffusion models, demonstrating significant advancements. By providing an all-in-one system, we offer a more efficient and effective pathway for community development in the field of image generation.
2 Related Work
2.1 Text-based Image Generation
Initially, Generative Adversarial Networks (GANs) [15, 26] were widely used as the primary approach for Text-based image generation. However, the landscape of image generation has evolved, and diffusion models [6] have emerged as a dominant framework, especially when integrated with text encoders such as CLIP [12] and T5 [13], enabling precise text-conditioned image generation. For instance, DAELL-2 [14] leverages CLIP’s image embeddings, derived from CLIP’s text embeddings through a prior model, to generate high-quality images. Similarly, Stable Diffusion [16] directly generates images from CLIP’s text embeddings. Imagen [17], on the other hand, utilizes a powerful language model like T5 [13] to encode text prompts, resulting in accurate image generation. Transformer-based architectures have also demonstrated their efficacy in generating images from textual inputs. CogView2 [4] and Muse [2] are notable examples of such models. To align text-to-image diffusion models with human preferences, recent methods [8, 1, 24] propose training diffusion models with reward signals. This ensures that the generated images not only meet quality benchmarks but also closely align with human intent and preferences. These cutting-edge approaches aim to improve the fidelity and relevance of generated images to better satisfy user requirements.
Figure 2: Overview of DiffusionGPT. The workflow of DiffusionGPT consists of four steps: Prompt Parse, Tree-of-thought of Models of Building and Searching, Model Selection, and Execution Generation. The four steps are shown from left to right and interact with LLM continuously. The upper side shows the detailed process of each step. The lower side shows the example of the whole workflow.
2.2 Large Language Models (LLMs) for Vision-Language Tasks
The field of natural language processing (NLP) has witnessed a significant transformation with the emergence of large language models (LLMs) [9, 3, 21], which have demonstrated remarkable proficiency in human interaction through conversational interfaces. To further enhance the capabilities of LLMs, the Chain-of-Thought (CoT) framework [22, 28, 7, 29] has been introduced. This framework guides LLMs to generate answers step-by-step, aiming for superior final answers. Recent research has explored innovative approaches by integrating external tools or models with LLMs [11, 18, 20, 23, 19]. For example, Toolformer [18] empowers LLMs with the ability to access external tools through API tags. Visual ChatGPT [23] and HuggingGPT [19] extend the capabilities of LLMs by enabling them to leverage other models to handle complex tasks that go beyond language boundaries. Similarly, Visual Programming [5] and ViperGPT [20] harness the potential of LLMs in processing visual objects by utilizing programming languages to parse visual queries. Drawing inspiration from these endeavors, we embrace the concept of LLMs as versatile tools and leverage this paradigm to guide T2I models to generate high-quality images.
3 Methodology
Figure 3: Details of prompts during interactions with the ChatGPT[9]. Before being inputted into the ChatGPT, the slots “{}” in figure are uniformly replaced with the corresponding text values.
@@ 3 Methodology
DiffusionGPT is an all-in-one system specifically designed to generate high-quality images for diverse input prompts. Its primary objective is to parse the input prompt and identify the generative model that produces the most optimal results, which is high-generalization, high-utility, and convenient.
@@
DiffusionGPT composes of a large language model (LLM) and various domain-expert generative models from the open-source communities (e.g. Hugging Face, Civitai). The LLM assumes the role of the core controller and maintains the whole workflow of the system, which consists of four steps: Prompt Parse, Tree-of-thought of Models of Building and Searching, Model Selection with Human Feedback, and Execution of Generation. The overall pipeline of DiffusionGPT is shown in Figure 2.
3.1 Prompt Parse
@@
By identifying these forms of prompts, Prompt Parse Agent allows DiffusionGPT to accurately recognize the core content that the user wants to generate, while mitigating the influence of noisy text. This process is crucial for selecting appropriate generative models and achieving high-quality generation results.
3.2 Tree-of-thought of Models
@@
Following the prompt parsing stage, the subsequent step involves selecting appropriate generative models from an extensive model library to generate the desired images. However, considering the large number of models available, it is impractical to input all models simultaneously into the large language model (LLM) for selection. Additionally, as different models may exhibit similarities in their generation spaces, accurately identifying the most suitable model through a single fuzzy match across the entire model library becomes challenging. To address this issue and identify the optimal model, we propose the utilization of a model tree based on the concept of Tree-of-Thought (TOT).By leveraging the search capabilities of the model tree, we can narrow down the candidate set of models and enhance the accuracy of the model selection process.
@@
Constructing the Model Tree using TOT. The Tree-of-Thought (TOT) of Model Building Agent is employed to automatically build the Model Tree based on the tag attributes of all models. By inputting the tag attributes of all models into the agent, it analyzes and summarizes potential categories derived from the Subject Domain and Style Domain. The Style categories are then incorporated as subcategories within the Subject category, establishing a two-layer hierarchical tree structure. Subsequently, all models are assigned to suitable leaf nodes based on their attributes, thereby completing the comprehensive model tree structure. The figure below illustrates the visual representation of the model tree. As the model tree is automatically constructed by the agent, this approach ensures convenient extensibility for incorporating new models. Whenever new models are added, agent seamlessly places them in the appropriate position within the model tree based on their attributes.
@@
Searching the Model Tree using TOT. The search process within the Model Tree, based on the Tree-of-Thought (TOT) of Models Searching Agent, aims to identify a candidate set of models that closely align with the given prompt. This search method employs a breadth-first approach, systematically evaluating the best subcategory at each leaf node. At each level, the categories are compared against the input prompt to determine the category that exhibits the closest match. This iterative process continues to derive the candidate set for the subsequent leaf node, and the search progresses until reaching the final node, where the candidate set of models is obtained. This candidate set of models serves as the basis for model selection in the subsequent stage.
3.3 Model Selection
@@
The model selection stage aims to identify the most suitable model for generating the desired image from the candidate set obtained in the previous stage. This candidate set represents a subset of the entire model library, consisting of models that exhibit relatively high matching degrees with the input prompt. However, the limited attribute information available from open-source communities poses challenges in precisely determining the best model while providing detailed model information to the large language model (LLM). To address this, we propose a Model Selection Agent that leverages human feedback and utilizes advantage database technology to align the model selection process with human preferences.
@@
For the advantage database, we employ a reward model to calculate scores for all model-generated results based on a corpus of 10,000 prompts, storing the score information. Upon receiving an input prompt, we calculate the semantic similarity between the input prompt and the 10,000 prompts, identifying the top 5 prompts with the highest similarity. Subsequently, the Model Selection Agent retrieves the precomputed performance of each model for these prompts from an offline database and selects the top 5 models for each selected prompt. This process produces a candidate set of 5x5 models.
The agent then intersects the model set with the model candidate set obtained in the TOT of models stage, focusing on models with higher occurrence probabilities and relatively higher rankings. These models are ultimately chosen as the final selection for model generation.
3.4 Execution of Generation
@@
Prompt Extension. To enhance the quality of prompts during the generation process, a Prompt Extension Agent is employed to augment the prompt. This agent leverages prompt examples from the selected model to automatically enrich the input prompt. The example prompts and the input prompts are both sent to LLM in the in-context learning paradigm. Specially, this agent incorporates rich descriptions and detailed vocabulary to the input prompts following the sentence pattern of example prompts. For example, if the input prompt is “an image of a laughing woman, fashion magazine cover” and example prompt is “fashion photography portrait of woman avatar, 1girl in blue lush Alien Rainforest with flowers and birds, fantasy, octane render, hdr, Dolby Vision, (intricate details, hyperdetailed:1.2), (natural skin texture, hyperrealism, soft light:1.2), fluffy short hair, sharp focus, night, necklace, Chinese mythology, cleavage, medium breasts, sci-fi headband, looking at viewer, best quality, perfect body”, the Prompt Extension Agent enhances it to a more detailed and expressive form such as: “The woman on the magazine cover is laughing joyfully, her eyes twinkling with delight. She is wearing a fashionable outfit that accentuates her curves, and her hair is styled in a way that complements her features”. This augmentation significantly improves the quality of the generated outputs.
4 Experiments
Figure 4: When comparing SD1.5-based DiffusionGPT with SD15[16], it is observed that DiffusionGPT excels in generating more realistic results at a fine-grained level for categories such as humans and scenes. The generated images demonstrate improved visual fidelity, capturing finer details and exhibiting a higher degree of realism compared to SD15.
4.1 Settings
In our experimental setup, the primary large language model (LLM) controller employed was ChatGPT[9], specifically utilizing the text-davinci-003 version, which is accessible through the OpenAI API. To facilitate the guidance of LLM responses, we adopted the LangChain framework, which effectively controlled and directed the generated outputs. For the generation models utilized in our experiments, we selected a diverse range of models sourced from the Civitai and Hugging Face communities. The selection process involved choosing the most popular models across different types or styles available on these platforms.
Table 1: Quantitative Results: We evaluate the aesthetic score with image-reward and aesthetic score. Compare to “SD15” and “Random” select expert models to output, both Tree-of-Thoughts and Human Feedback can improve the aesthetic of generated images and human alignment.
@@ 4.1 Settings
DiffusionGPT wo HF 0.56 5.62
DiffusionGPT 0.63 5.70
4.2 Qualitative Results
4.2.1 Visualization of SD1.5 Version
@@ 4.2 Qualitative Results
To assess the efficacy of our system, we performed a comprehensive evaluation by comparing its generation performance against the baseline method, SD 1.5[16]. SD 1.5 serves as the foundational model for various specialized community models. The comparative results are presented in Figure 4. We conducted a detail analysis of four distinct prompt types and compared them along two key dimensions: semantic alignment and image aesthetics.
@@
Figure 5: Comparison of SDXL version of DiffusionGPT with baseline SDXL[10]. All generated iamges are $\times×1024$ pixels.
Figure 6: Comparison of DiffusionGPT-Xl with base model.
Figure 7: User Study: Comparing DiffusionGPT with SD1.5. Users strongly prefer expert models selected by DiffusionGPT over the baseline in terms of prompts from all 10 categories.
@@
Upon careful examination of the results, we have identified two notable issues with the base model: i) Semantic Lack: The base model’s generated images exhibit a limited focus on specific semantic classes derived from the input prompt, resulting in an incomplete capture of the overall semantic information. This limitation is particularly evident in all kinds of prompt types, where the base model struggles to effectively generate objects related to “man, chef, children, and toyland”. ii) Poor performance on human-related targets: The base model faces challenges in generating accurate facial and body details for human-related objects, leading to subpar image quality. This deficiency becomes apparent when comparing the aesthetic qualities of images depicting “girl and couple”.
@@
In contrast, DiffusionGPT effectively addresses these limitations. The images produced by our system demonstrate relatively complete representations of the target regions, successfully capturing the semantic information encompassing the entire input prompt. Examples such as “man who whistles tunes pianos” and “a snowy wonderland where children build snowmen and have snowball fights” showcase the capacity of our system to encompass the broader context. Furthermore, our system excels in generating more detailed and accurate image for human-related objects. This is exemplified by the prompt “a romantic couple sharing a tender moment under a starry sky”.
@@
Figure 8: Ablation study of DiffusionGPT. The random selection is the baseline method for generating images. The TOT or TOT+HF represent the performance of different agents.
@@
Figure 9: Ablation study of Prompt Extension. The extension aims to provide the riched prompts that produces higher quality images.
4.2.2 Visualization of SDXL Version.
With the advancements in publicly available universal generation models, the newly improved method SD XL[10] has emerged as a promising approach, demonstrating superior generation results. To further enhance our system, we have developed an upgraded version by integrating various open-source community models based on SD XL. In order to evaluate the performance of our system, we compared it against SD XL, as depicted in Figure 5. It is important to note that all output images have a resolution of 1024x1024, and four distinct types of prompts were generated for the purpose of comparison. Upon careful analysis, it becomes apparent that SD XL occasionally exhibits a partial loss of semantic information in specific cases. For instance, the generated results for prompts involving “3D tiger” in the prompt-based category or “flying cars” in the instruction-based category may lack accurate representation. In contrast, our system excels in producing more precise and visually appealing representations. Noteworthy examples include the generation of “a white towel with a cartoon cat” and the depiction of a “starry sky”.
4.3 Quantitative Results

### v2

# DiffusionAgent[Uncaptioned image]: Navigating Expert Models for Agentic Image Generation

@@ Abstract
In the accelerating era of human-instructed visual content creation, diffusion models have demonstrated remarkable generative potential. Yet their deployment is constrained by a dual bottleneck: semantic ambiguity in diverse prompts and the narrow specialization of individual models. A single diffusion architecture struggles to maintain optimal performance across heterogeneous prompts, while conventional “parse-then-call” pipelines artificially separate semantic understanding from generative execution. To bridge this gap, we introduce DiffusionAgent, a unified, language-model-driven agent that casts the entire “prompt comprehension–expert routing–image synthesis” loop into a agentic framework. Our contributions are three-fold: (1) a tree-of-thought-powered expert navigator that performs fine-grained semantic parsing and zero-shot matching to the most suitable diffusion model via an extensible prior-knowledge tree; (2) an advantage database updated with human-in-the-loop feedback, continually aligning model-selection policy with human aesthetic and semantic preferences; and (3) a fully decoupled agent architecture that activates the optimal generative path for open-domain prompts without retraining or fine-tuning any expert. Extensive experiments show that DiffusionAgent retains high generation quality while significantly broadening prompt coverage, establishing a new performance and generality benchmark for multi-domain image synthesis. The code is available at https://github.com/DiffusionAgent/DiffusionAgent.
@@
Figure 1: We propose a unified generation system DiffusionAgent, which leverages Large Language Models (LLM) to seamlessly accommodate various types of prompts input and integrate domain-expert models for output. Our system is capable of parsing diverse forms of inputs, including Prompt-based, Instruction-based, Inspiration-based, and Hypothesis-based input types. It exhibits the ability to generate outputs of superior quality.
I Introduction
@@
Recent years have witnessed the prevalence of diffusion models[5, 15, 11, 16] in human-interactive image generation tasks, revolutionizing image editing, stylization, and other related tasks. The first open-source text-to-image diffusion model, known as Stable Diffusion (SD)[15], which has rapidly gained popularity and widespread usage. Various techniques tailored for SD, such as Controlnet[27], Lora[6], further paved the way for the development of SD and foster its integration into various applications. SDXL[11] is tailored to deliver exceptional photorealistic outputs with intricate details and artistic composition. Moreover, the impact of SD extends beyond technical aspects. Community platforms such as Civitai, WebUI, and LibLibAI have emerged as vibrant hubs for discussions and collaborations among designers and creators.
@@
Despite making significant strides, current stable diffusion models face two key challenges when applied to realistic scenarios as shown in Fig. 1:
  • Model Limitation: Stable diffusion models like SD1.5[15] are versatile but underperform in specific domains, whereas domain-specific models like SD1.5+Lora excel in targeted sub-fields at the cost of general adaptability.
@@
  • Prompt Constraint: During stable diffusion training, text data only includes descriptive captions, yet the models face challenges in achieving optimal performance with diverse prompt, such as instructions and inspirations.
Figure 2: Overview of DiffusionAgent. The workflow of DiffusionAgent consists of four steps: Prompt Parse, Tree-of-thought of Models of Building and Searching, Model Selection, and Execution Generation. The four steps are shown from left to right and interact with LLM continuously. The upper side shows the detailed process of each step. The lower side shows the example of the whole workflow.
@@
The mismatch between stable diffusion models and real-world applications often results in limited performance, poor generalization, and implementation difficulties. Research has sought to address these issues. While SDXL has improved specific-domain performance, but optimal results are still unattained. Techniques like prompt engineering and fixed templates have been used to enhance input quality and output, yet they fall short of offering a complete solution. This leads us to pose a key question: Can we create a unified framework to unleash prompt constraint and activate corresponding domain expert model ?
@@
In order to address the aforementioned question, we propose DiffusionAgent, which leverages Large Language Model (LLM)  [10, 21, 1] to offer a one-for-all generation system that seamlessly integrates superior generative models and effectively parses diverse prompts. DiffusionAgent constructs a Tree-of-Thought (ToT) structure, which encompasses various generative models based on prior knowledge and human feedback. When presented with an input prompt, the LLM first parses the prompt and then guides the ToT to identify the most suitable model for generating the desired output. Furthermore, we introduce Advantage Databases, where the Tree-of-Thought is enriched with valuable human feedback, aligning the LLM’s model selection with human preferences.
@@
  • New Insight: DiffusionAgent utilizes a Large Language Model (LLM) as the cognitive engine for its text-to-image generation system, processing various inputs and enabling expert selection for outputs.
@@
  • All-in-one System: DiffusionAgent provides a versatile solution compatible with various diffusion models, unlike existing approaches that are restricted to descriptive prompts, thereby broadening its applicability to different prompt types.
@@
  • Efficiency and Pioneering: DiffusionAgent is notable for its training-free, plug-and-play integration. By incorporating the Tree-of-Thought (ToT) and human feedback, our system enhances accuracy and facilitates a flexible process for aggregating multiple experts.
@@
  • High Effectiveness: DiffusionAgent outperforms traditional stable diffusion models, showcasing significant advancements and offering an all-in-one system that provides a more efficient and effective pathway for community development in image generation.
II Methodology
@@
DiffusionAgent is an all-in-one system specifically designed to generate high-quality images for diverse input prompts. Its primary objective is to parse the input prompt and identify the generative model that produces the most optimal results, which is high-generalization, high-utility, and convenient.
@@
DiffusionAgent composes of a large language model (LLM) and various domain-expert generative models from the open-source communities (e.g. Hugging Face, Civitai). The LLM acts as the main controller and maintains the workflow of the system, which consists of four steps: Prompt Parse, Tree-of-thought of Models of Building and Searching, Model Selection with Human Feedback, and Execution of Generation. The overall pipeline of DiffusionAgent is shown in Fig. 2.
II-A Prompt Parse
@@
The Prompt Parse Agent helps DiffusionAgent identify core content from prompts, reducing the impact of noisy text. This is essential for selecting suitable generative models and achieving high-quality results.
II-B Tree-of-thought of Models
@@
Following the prompt parsing stage, the subsequent step involves selecting appropriate generative models from an extensive model library to generate the desired images. However, considering the large number of models available, it is impractical to input all models simultaneously into the large language model (LLM) for selection. To address this issue and pinpoint the optimal model, we draw inspiration from the Chain-of-Thought (CoT) paradigm [22, 28] and introduce a Tree-of-Thought (ToT) model tree. By leveraging the search capabilities of the model tree, we can narrow down the candidate set of models and enhance the accuracy of the model selection process.
@@
Constructing the Model Tree using TOT. The Tree-of-Thought (TOT) of Model Building Agent is employed to automatically build the Model Tree based on the tag attributes of all models. The agent processes the tag attributes of all models to analyze and categorize them into Subject and Style Domains. Style categories become subcategories within the Subject category, forming a two-layer hierarchical tree. Models are assigned to appropriate leaf nodes based on their attributes, completing the model tree. This automatically constructed tree allows for easy extensibility, as new models are seamlessly integrated based on their attributes.
@@
Searching the Model Tree using TOT. The Model Tree search process, guided by the Tree-of-Thought (TOT) of Models Searching Agent, aims to find models closely aligned with a given prompt. It uses a breadth-first search, systematically evaluating the best subcategory at each leaf node. Categories are compared against the prompt to find the closest match. This iterative process continues until the final node, yielding a candidate set of models for next selection.
II-C Model Selection
@@
The model selection stage aims to identify the most suitable model for generating the desired image from the candidate set obtained in the previous stage. The candidate set, a subset of the model library, includes models closely matching the input prompt. However, limited attribute data from open-source sources complicates precise model selection and detailed information provision to the LLM. To tackle this, we propose a Model Selection Agent that uses human feedback and advanced database technology to align model selection with human preferences.
@@
The advantage database uses a reward model to score model-generated results from 10,000 prompts, storing these scores. For a given input prompt, we calculate its semantic similarity to these prompts and identify the top 5 matches. The Model Selection Agent then retrieves precomputed model performances for these prompts from an offline database, selecting the top 5 models for each. This results in a candidate set of 25 models. The agent intersects the model set with the candidate set from the TOT stage, prioritizing models with higher occurrence probabilities and rankings. These models are chosen for the final model generation.
Figure 3: Details of prompts during interactions with the LLM. Before being inputted into the LLM, the slots “{}” in figure are uniformly replaced with the corresponding text values.
II-D Execution of Generation
@@
Prompt Extension. To enhance the quality of prompts during the generation process, a Prompt Extension Agent is employed to augment the prompt. This agent leverages prompt examples from the selected model to automatically enrich the input prompt. The example prompts and the input prompts are both sent to LLM in the in-context learning paradigm. Specially, this agent incorporates rich descriptions and detailed vocabulary to the input prompts following the sentence pattern of example prompts. The Prompt Extension Agent enhances it to a more detailed and expressive form, which significantly improves the quality of the outputs.
Figure 4: When comparing SD1.5-based DiffusionAgent with SD1.5[15], it is observed that DiffusionAgent excels in generating more realistic results at a fine-grained level for categories such as humans and scenes. The generated images demonstrate improved visual fidelity, capturing finer details and exhibiting a higher degree of realism compared to SD1.5.
III Experiments
III-A Settings
In our experimental setup, the primary large language model (LLM) controller employed was ChatGPT[10], specifically utilizing the text-davinci-003 version, which is accessible through the OpenAI API. To facilitate the guidance of LLM responses, we adopted the LangChain framework, which effectively controlled and directed the generated outputs. For the generation models utilized in our experiments, we selected a diverse range of models sourced from the Civitai and Hugging Face communities. The selection process involved choosing the most popular models across different types or styles available on these platforms. The selected models for different versions of DiffusionAgent are illustrated in Table I. The details of prompts during interactions with LLM are shown in Fig. 3.
TABLE I: The selected public trained models for SD1.5 and SDXL based system.
SD1.5 Based SDXL Based
FilmVelvia2 SDXL
majicmixRealistic weird-future-fashion
lou fenrisxl-801Photorealistic
cartoonish dynavisionXLAllInOneStylized
animemix sdxlNijiSpecial
impasto-painting copaxTimelessxlSDXL1
CineStyle5 starcitizen
dreamlabsoil lwmirXL
ghibli-style kandinsky
dreamshaper dreamshaperXL
C4D
TABLE II: Quantitative Results: We evaluate the aesthetic score using image-reward and aesthetic score compare to “SD1.5” and “Random” select expert models to output.
@@ III-A Settings
DiffusionAgent wo HF 0.56 5.62
DiffusionAgent 0.63 5.70
III-B Qualitative Results
III-B1 Visualization of SD1.5 Version
@@ III-B Qualitative Results
To assess the efficacy of our system, we performed a comprehensive evaluation by comparing its generation performance against the baseline method, SD1.5[15]. The comparative results are presented in Fig. 4. We conducted a detail analysis of four distinct prompt types and compared them along two key dimensions: semantic alignment and image aesthetics.
@@
Figure 5: Comparison of SDXL version of DiffusionAgent with baseline SDXL[11]. All generated iamges are $\times1024$ pixels.
@@
Upon examining the results, two issues with the base model were identified: i) Semantic Lack: The base model struggles to capture complete semantic information, particularly with prompts involving “man, chef, children, and toyland”, focusing too narrowly on specific classes. ii) Poor Performance on Human-Related Targets: The base model has difficulty generating accurate facial and body details, resulting in lower quality images, especially for “girl and couple” prompts.
@@
In contrast, DiffusionAgent effectively addresses these limitations by producing images that capture complete semantic information from input prompts. For example, it successfully represents broader contexts like “man who whistles tunes pianos” and “a snowy wonderland where children build snowmen”. Additionally, it excels in generating detailed images of human-related objects, as seen with prompts like “a romantic couple sharing a tender moment under a starry sky”.
@@
Figure 6: Ablation study of DiffusionAgent. The random selection is the baseline method for generating images. The TOT or TOT+HF represent the performance of different agents.
@@
Figure 7: Ablation of Prompt Extension. It aims to provide the riched prompts that produces higher quality images.
III-B2 Visualization of SDXL Version.
With advancements in universal generation models, SDXL[11] has shown promising results. We improved our system by integrating various open-source models based on SDXL. All comparative results are shown in Fig. 5, all output images were 1024x1024. The result reveals that SDXL sometimes loses semantic information, as seen in prompts like “3D tiger” or “flying cars”. In contrast, our system produces more visually appealing images, such as “a white towel with a cartoon cat” and a “starry sky”.
III-C Quantitative Results